# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora. However, this poses risk of privacy and copyright violations, highlighting the need for efficient machine unlearning methods that remove sensitive data without retraining from scratch. While Gradient Ascent (GA) is commonly used to unlearn by reducing the likelihood of generating unwanted content, it leads to unstable optimization and catastrophic forgetting of retrained knowledge. We also find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance. To address these challenges, we propose two novel techniques for robust and efficient unlearning for LLMs. First, we introduce Inverted Hinge loss, which suppresses unwanted tokens while maintaining fluency by boosting the probability of the next most likely token. Second, we develop a data-adaptive initialization for LoRA adapters via low-rank approximation weighted with relative Fisher information, thereby focusing updates on parameters critical for removing targeted knowledge. Experiments on the Training Data Extraction Challenge dataset using GPT-Neo models as well as on the TOFU benchmark with Phi-1.5B and Llama2-7B models demonstrate that our approach effectively removes sensitive information while maintaining reasoning and generative capabilities with minimal impact.
@@ 1 Introduction
One method for LLM unlearning would be to filter out sensitive data from the corpus and retrain the model from scratch, an approach known as exact unlearning. With unprecedentedly large models and pretraining datasets, this process is highly resource-intensive and can easily become intractable under the possibility of multiple data deletion requests made in a sequential manner. This motivates approximate unlearning, where the goal is to remove knowledge of specific data instances without retraining the model from scratch (Figure 1). In this regard, Jang et al. (2023) proposed a simple method that finetunes LLMs using Gradient Ascent (GA) on data requested for deletion, and also introduced $n$-gram-based metrics to assess its efficacy. However, the proposed GA method suffers significantly not only from unstable optimization due to the objective loss being unbounded, but also from high computational cost of full-finetuning all parameters within the LLM.
Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it forgets knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion. After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other relevant data points in the retain set ${\mathcal{D}}_{r}$ as well as reasoning and generative capabilities acquired from pretraining.
@@
Meanwhile, Low-Rank Adaptation (LoRA) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning on downstream tasks (Hu et al., 2022). The core idea of LoRA is to freeze all pretrained weights and instead train low-rank decomposition matrices to model the weight changes in each linear layer, effectively reducing the number of trainable parameters and thus its memory cost. In addition to its efficiency, the low-rankness in LoRA also induces a powerful regularization (Biderman et al., 2024), which can be beneficial in the realm of LLM unlearning by stabilizing optimization and preventing catastrophic forgetting of other remaining knowledge. From this perspective, we conjecture that LoRA would be a valuable approach to consider in practical unlearning scenarios, yet its application to LLM unlearning remains unexplored.
@@
In this paper, we explore LLM unlearning under the low-rank adaptation paradigm and propose two novel techniques for robust and efficient knowledge unlearning for LLMs. First, we analyze the derivatives of GA and highlight its shortcomings: 1) gradients increasing the probability of all other possible tokens cause unnecessary forgetting and 2) maximizing the next-token prediction loss involves unbounded optimization and can easily diverge. To address these issues, we propose the Inverted Hinge Loss (IHL) that aims to replace each token to unlearn with the next most-probable token, and show that IHL enables fast and stable tuning by resolving the issues of GA. Second, we find that the low-rank of LoRA imposes regularization too strong for LLM unlearning, and forcing the model to fully unlearn data points through extensive tuning leads to suboptimal cost vs. post-unlearning performance trade-offs. To accelerate unlearning in a controlled manner, we propose Fisher-Initialization of Low-rank Adapters (FILA), which data-adaptively assigns parameters responsible for generating unwanted information to adapters prior to tuning by decomposing the pretrained parameters weighted by the relative Fisher-information matrix. Our extensive experiments on unlearning samples in the Training Data Extraction Challenge dataset with GPT-Neo models as well as removing knowledge of fictitious author profiles using the TOFU benchmark with Phi-1.5B and Llama2-7B models consistently demonstrate that our proposed IHL combined with FILA significantly outperforms previous baselines in both efficiency and post-unlearning downstream performance.
@@
Figure 2: Compute cost for successful unlearning vs. post-unlearning downstream performances. We unlearn 32 randomly sampled sequences from the Training Data Extraction Challenge from GPT-Neo-125M. Each point represents a different forget set and LoRA rank (if used). Left: Accuracy averaged across 9 classification tasks (higher is better). Middle: F1 score averaged across 4 dialogue generation tasks (higher is better). Right: Perplexity on the validation set of the Pile dataset (lower is better). Dashed lines indicate the performances of the model prior to unlearning. Unlearning via gradient differences (GD) with vanilla LoRA leads to significant loss in performance compared to full-parameter GD unlearning due to lack of plasticity. However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the cost-efficiency of LoRA.
@@ 2 Related Work
Machine Unlearning. The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pretrained model to discard information acquired from a specific subset of data previously used for pretraining. Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2023; Mehta et al., 2022; Chundawat et al., 2023; Cha et al., 2024). Recently, its significance has grown notably with Large Language Models (LLMs) due to crucial need for managing unintended memorization of pretraining data intrinsic to LLMs (Si et al., 2023; Yao et al., 2024b).
Several machine unlearning algorithms have been proposed for LLMs, with most falling under the category of parameter optimization-based methods (Si et al., 2023). First, Wang et al. (2023) introduced the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets. Chen & Yang (2023) proposed the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model intact. Liu et al. (2024) developed a two-stage framework: the harmful knowledge acquisition stage identifies and captures harmful knowledge within the model, and the knowledge negation stage removes that knowledge. While these approaches have demonstrated impressive unlearning results across various datasets, they are limited in applicability either by the need to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) or by the need to infer from a secondary model for knowledge distillation (Wang et al., 2023; Liu et al., 2024). Another category of LLM unlearning methods are tuning-free model-editing methods, an example of which leverages task arithmetic to suppress harmful text (Ilharco et al., 2023; Wu et al., 2023). While free from cost of tuning, these methods fails to show sufficient level of unlearning (Ilharco et al., 2023) or incur higher computational costs for detecting privacy neurons (Wu et al., 2023).
In contrast, inspired by image classifier unlearning literature (Golatkar et al., 2020; Cha et al., 2024), Jang et al. (2023) adopted Gradient Ascent (GA) to unlearn from LLMs by maximizing the next-token prediction loss for all sequences in the forget set. Unlike in image classification, they demonstrate that LLMs show negligible forgetting of retained knowledge when using GA on the forget data. Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks. More recently, Yao et al. (2024a) demonstrated that incorporating GA with gradient descent on in-distribution data further enhances robustness against hyperparameters.
@@
Parameter-Efficient Fine-Tuning. Full-finetuning LLMs to adapt towards specific downstream tasks and instructions incurs an intractable computational cost due to the large model size and complexity. To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023). Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2022; Zhang et al., 2023; Yeh et al., 2023; Kopiczko et al., 2024; yang Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter. A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM. While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposed to initialize LoRA weights using the principal singular vectors and values of the linear weights.
@@ 3 Proposed Method
In this section, we introduce useful notation and provide background information on unlearning with Gradient Ascent (GA) and Low-Rank Adaptation (LoRA). We then analyze derivatives of GA and highlight its inherent issues. Building upon this analysis, we present the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA. Lastly, we illustrate the method of initializing LoRA data-adaptively via Fisher-information, thereby enabling cost-efficient unlearning.
@@ 3.1 Preliminaries
Gradient Ascent. Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, leading to a simple yet effective baseline of Gradient Ascent (GA; Jang et al. 2023). GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
$\mathcal{L}_{\text{GA}}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t})).$ (1)
In practice, the log-likelihood is computed using cross-entropy loss and thus we GA essentially minimizes the Negative Cross-Entropy (NCE) loss. Therefore, GA maximizing the next-token prediction loss involves unbounded optimization, leading to an ill-posed process with unstable tuning. While Gradient Difference (GD) aims to alleviate this instability by minimizing the next-token prediction loss for ${\mathcal{D}}_{r}$ alongside NCE on ${\mathcal{D}}_{f}$ as regularization, we find that the approach falls short of a fundamental solution, showing performance degradation as unlearning updates are made.
Low-Rank Adaptation. Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank (Aghajanyan et al., 2021), LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter. In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes:
$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}.$
@@ 3.3 Inverted Hinge Loss: A Novel Loss Function for LLM Unlearning
From this derivative of GA, we can interpret its unlearning mechanism: GA decreases the prediction score corresponding to the true token $x_{t}$, while increasing the prediction score for all other tokens $v\neq x_{t}$ in the vocabulary. Consequently, we hypothesize that GA suffers from the following problems: 1) the objective is unbounded with the possibility of divergence during optimization, 2) gradient updates are inefficient due to gradients spreading across all other logits with large vocabulary sizes, and 3) there is unnecessary degradation of generative performance due to increasing the logits for all other tokens regardless of textual coherence.
Inverted Hinge Loss. To cope with aforementioned limitations of GA, we aim to design a new loss function that achieves effective unlearning by decreasing the prediction score of the true token, while focusing gradient updates on only a minimal number of viable replacements for the ground-truth token. Inspired by the Hinge loss (Cortes & Vapnik, 1995), we devise Inverted Hinge loss (IHL) as below:
$\mathcal{L}_{\text{IHL}}({\bm{x}})=\max\left(0,1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))\right)$
@@
$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}}.\end{cases}$
Here, $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$ and the above represents the derivative when $\mathcal{L}_{\text{IHL}}({\bm{x}})\neq 0$. Note that $\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=0$ when $\mathcal{L}_{\text{IHL}}({\bm{x}})=0$. The detailed derivation can be found in Appendix A.
The derivative of $p_{\theta}(x_{t}|x_{<t})$ and $p_{\theta}(v^{\star}|x_{<t})$ clearly illustrate how the IHL addresses the shortcomings of GA in knowledge unlearning for LLMs. First, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})$ is greater than $p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with opposite sign). This ensures that not only knowledge unlearning for the $t$-th token is executed rapidly but also prevents spreading out of gradients. During this process, the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ increase slowly in proportion to the difference between the predictions for $x_{t}$ and $v^{\star}$. Second, once knowledge unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$), the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease. This not only prevents unnecessary forgetting but also results in a bounded loss function, allowing stable unlearning via stochastic gradient descent.

### v3

# Towards Robust and Parameter-Efficient Knowledge Unlearning for LLMs

@@ Abstract
Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora. However, this poses risk of privacy and copyright violations, highlighting the need for efficient machine unlearning methods that remove sensitive data without retraining from scratch. While Gradient Ascent (GA) is commonly used to unlearn by reducing the likelihood of generating unwanted content, it leads to unstable optimization and catastrophic forgetting of retrained knowledge. We find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance. To address these challenges, we propose two novel techniques for robust and efficient unlearning for LLMs. First, we introduce Inverted Hinge Loss, which suppresses unwanted tokens while maintaining fluency by boosting the probability of the next most likely token. Second, we develop a data-adaptive initialization for LoRA adapters via low-rank approximation weighted with relative Fisher information, thereby focusing updates on parameters critical for removing targeted knowledge. Experiments on the Training Data Extraction Challenge dataset using GPT-Neo models as well as on the TOFU benchmark with Phi-1.5B and Llama2-7B models demonstrate that our approach effectively removes sensitive information while maintaining reasoning and generative capabilities with minimal impact. Our implementation can be found in https://github.com/csm9493/efficient-llm-unlearning.
@@ 1 Introduction
One method for LLM unlearning would be to filter out sensitive data from the corpus and retrain the model from scratch, an approach known as exact unlearning. With unprecedentedly large models and pretraining datasets, this process is highly resource-intensive and can easily become intractable under the possibility of multiple data deletion requests made in a sequential manner. This motivates approximate unlearning, where the goal is to remove knowledge of specific data instances without retraining the model from scratch (Figure 1). In this regard, several novel approaches have been proposed for approximate unlearning: Jang et al. (2023) introduced a simple method that finetunes LLMs using Gradient Ascent (GA) on data requested for deletion and also proposed $n$-gram-based metrics to evaluate its effectiveness. Wang et al. (2023) and Liu et al. (2024) proposed knowledge distillation-based methods that selectively transfer knowledge to a secondary model for unlearning. However, both approaches face significant challenges: GA suffers from unstable optimization due to unbounded nature of its objective loss, while distillation-based methods incur substantial computational costs from relying on a secondary model. Above all, these approaches share a critical drawback: the high computational cost of full fine-tuning all parameters within the LLMs.
Figure 1: LLM unlearning aims to forget data points in ${\mathcal{D}}_{f}$ while maintaining knowledge of the retain set ${\mathcal{D}}_{r}$. Unlike GA, our IHL induces higher unlearning stability by reducing the likelihood of unwanted tokens in a controlled manner. To accelerate unlearning with IHL, FILA extracts and places parameters important in generating $\mathcal{D}_{f}$ to LoRA weights a priori via weighted low-rank approximation. IHL and FILA form a powerful synergy towards robust and efficient LLM unlearning.
@@
Meanwhile, Low-Rank Adaptation (LoRA) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning on downstream tasks (Hu et al., 2022). The core idea of LoRA is to freeze all pretrained weights and instead train low-rank decomposition matrices to model the weight changes in each linear layer, effectively reducing the number of trainable parameters and thus its memory cost. Beyond its efficiency, the low-rank structure in LoRA also serves as a strong regularizer (Biderman et al., 2024), which we hypothesize aids LLM unlearning by stabilizing optimization and mitigating catastrophic forgetting of retained knowledge. However, the empirical effects of LoRA in the context of LLM unlearning remain largely unexplored.
@@
In this paper, we present the first in-depth study of LLM unlearning under the low-rank adaptation paradigm and introduce two novel techniques for robust and parameter-efficient knowledge unlearning. First, we analyze the derivatives of GA to highlight its drawbacks such as (1) unnecessary forgetting induced by indiscriminately increasing the probability of all tokens and (2) unstable optimization due to the unboundedness of maximizing the next-token prediction loss. To tackle these challenges, we introduce the Inverted Hinge Loss (IHL), which replaces each unwanted token with the next most-probable one, and demonstrate that IHL facilitates stable unlearning by resolving the limitations of GA. Second, we find that the low-rank regularization in LoRA is too strong when unlearning with IHL, leading to suboptimal cost vs. post-unlearning performance trade-offs. In response, we propose Fisher-Initialization of Low-rank Adapters (FILA), which data-adaptively assigns parameters responsible for generating unwanted information to adapters prior to tuning by decomposing the pretrained parameters weighted by the relative Fisher-information matrix (Figure 1). Experiments on the Training Data Extraction Challenge dataset (GPT-Neo) and the TOFU benchmark (Phi-1.5B, Llama2-7B) demonstrate that IHL combined with FILA outperforms existing baselines in both efficiency and post-unlearning performance. In summary, our main contributions are as follows:
  • We analyze the shortcomings of GA—unbounded optimization and unnecessary forgetting—through its derivative and propose IHL to address these issues.
  • We introduce FILA, a method to accelerate unlearning by data-adaptively assigning parameters responsible for unwanted information to low-rank adapters before tuning.
  • We demonstrate that IHL combined with FILA outperforms previous baselines in terms of both efficiency and post-unlearning performance (Figure 2).
@@
Figure 2: Compute cost for successful unlearning vs. post-unlearning downstream performances. We unlearn 32 randomly sampled sequences in the Training Data Extraction Challenge from GPT-Neo-125M. Each point represents a different forget set and LoRA rank (if used). Left: Accuracy averaged across 9 classification tasks (higher is better). Middle: F1 score averaged across 4 dialogue generation tasks (higher is better). Right: Perplexity on the validation set of the Pile dataset (lower is better). Dashed lines indicate the performances of the model prior to unlearning. Unlearning via gradient differences (GD) with vanilla LoRA leads to significant loss in performance compared to full-parameter GD unlearning due to lack of plasticity. However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the parameter-efficiency of LoRA.
@@ 2 Related Work
Machine Unlearning.   The primary objective of machine unlearning is to adapt a pretrained model to discard information acquired from a specific subset of data previously used during pretraining, with active research focused on image classification (Cao & Yang, 2015; Golatkar et al., 2020; Tarun et al., 2023; Mehta et al., 2022; Chundawat et al., 2023; Cha et al., 2024). Recently, its significance has grown notably with Large Language Models (LLMs) due to crucial need for managing unintended memorization of pretraining data intrinsic to LLMs (Si et al., 2023; Yao et al., 2024b). Several machine unlearning algorithms for LLMs focus on parameter optimization (Si et al., 2023). For example, Wang et al. (2023) introduced Knowledge Gap Alignment, using knowledge distillation between models trained on different datasets. Chen & Yang (2023) proposed an unlearning layer to selectively remove specific knowledge while preserving other parameters, while Liu et al. (2024) developed a two-stage framework to capture and negate harmful knowledge. However, these approaches are limited by the need to retain large datasets (Wang et al., 2023; Chen & Yang, 2023) or rely on secondary models for distillation (Wang et al., 2023; Liu et al., 2024). Model-editing methods, such as task arithmetic for suppressing harmful content (Ilharco et al., 2023; Wu et al., 2023), avoid substantial costs but show limited effectiveness in unlearning. In contrast, Jang et al. (2023) used Gradient Ascent (GA) for LLM unlearning by maximizing the next-token prediction loss on the forget data, effectively unlearning while preserving model performance. Following this, GA has become a standard baseline, with Yao et al. (2024a) improving its robustness by combining GA with gradient descent on in-distribution data.
@@
Parameter-Efficient Fine-Tuning.   Fine-tuning LLMs for specific tasks is computationally expensive due to their large parameter sizes. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods adapt only a small subset of parameters while keeping the pretrained ones frozen (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023). Inspired by the small intrinsic rank of LLMs (Li et al., 2018; Aghajanyan et al., 2021), LoRA and its derivatives add low-rank adapters to the model’s linear layers (Hu et al., 2022; Zhang et al., 2023; Yeh et al., 2023; Kopiczko et al., 2024; yang Liu et al., 2024). These adapters can be merged with pretrained parameters after fine-tuning, maintaining the original inference cost. While most methods use random initialization for LoRA adapters, PiSSA (Meng et al., 2024) suggests initializing them using the principal singular vectors and values of the linear weights.
Although various methods in machine unlearning and parameter-efficient fine-tuning for LLMs have been discussed, this paper focuses on analyzing the inherent issues of GA, introducing a novel unlearning loss function to overcome these issues, and exploring parameter-efficient unlearning methods that do not require full fine-tuning. These areas have not been adequately addressed in previous studies, highlighting the contributions of our work.
@@ 3.1 Preliminaries
Gradient Ascent. Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, leading to a simple yet effective baseline of Gradient Ascent (GA; Jang et al. 2023). GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss
$\mathcal{L}_{\text{GA}}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$ (1)
In practice, the log-likelihood is computed using cross-entropy loss and thus GA essentially minimizes the Negative Cross-Entropy (NCE) loss. Therefore, GA maximizing the next-token prediction loss involves unbounded optimization, leading to an ill-posed process with unstable tuning. While Gradient Difference (GD) aims to alleviate this instability by minimizing the next-token prediction loss for ${\mathcal{D}}_{r}$ alongside NCE on ${\mathcal{D}}_{f}$ as regularization, we find that the approach falls short of a fundamental solution, showing performance degradation as unlearning updates are made.
Low-Rank Adaptation. Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank (Aghajanyan et al., 2021), LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter. In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes
$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}$
@@ 3.3 Inverted Hinge Loss: A Novel Loss Function for LLM Unlearning
From this derivative of GA, we can interpret its unlearning mechanism: given the prefix $x_{<t}$, GA reduces the prediction score of the true token $x_{t}$ in proportion to $1-p_{\theta}(x_{t}|x_{<t})$ while increasing the scores of other tokens (i.e., $v\neq x_{t}$) by $p_{\theta}(v|x_{<t})$. This process effectively shifts the model’s prediction for $x_{<t}$ away from the true token $x_{t}$, thereby achieving unlearning. However, we can confirm that GA suffers form the following problems during unlearning: (1) Gradient spread, where reducing the score of $x_{t}$ while increasing the scores of all other tokens leads to inefficient unlearning in large vocabularies by predominantly boosting other tokens; (2) Unbounded loss, where minimizing $\log(p_{\theta}(x_{t}|x_{<t}))$ through maximizing cross-entropy loss introduces a risk of divergence due to the unbounded nature of entropy; and (3) Degradation of generative performance, where GA applies uniform gradient updates (i.e. increasing the scores of other tokens) to all sequences in the forget set $\mathcal{D}_{f}$, despite each sequence requiring a unique number of updates for unlearning. This redundancy can cause degrade the model’s generative capabilities, resulting in catastrophic forgetting.
Inverted Hinge Loss. To cope with aforementioned limitations of GA, we aim to design a new loss function that achieves effective unlearning by decreasing the prediction score of the true token, while focusing gradient updates on only a minimal number of viable replacements for the ground-truth token. Inspired by the Hinge Loss (Cortes & Vapnik, 1995), we devise Inverted Hinge Loss (IHL) as
$\mathcal{L}_{\text{IHL}}({\bm{x}})=1+p_{\theta}(x_{t}|x_{<t})-\max_{v\neq x_{t}}(p_{\theta}(v|x_{<t}))$
@@
$\frac{\partial\mathcal{L}_{\text{IHL}}({\bm{x}})}{\partial y_{t}^{(v)}}=\begin{cases}p_{\theta}(x_{t}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})+1)&\text{if }v=x_{t}\\ p_{\theta}(v^{\star}|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t})-1)&\text{if }v=v^{\star}\\ p_{\theta}(v|x_{<t})(p_{\theta}(v^{\star}|x_{<t})-p_{\theta}(x_{t}|x_{<t}))&\text{if}{\ v\neq x_{t}\ \text{and}\ v\neq v^{\star}},\end{cases}$
where $v^{\star}=\arg\max_{v\neq x_{t}}p_{\theta}(v|x_{<t})$. The detailed derivation can be found in Appendix A.
The above derivative clearly illustrates how the IHL addresses the shortcomings of GA in knowledge unlearning for LLMs. First, IHL mitigates gradient spread by ensuring that the gradients primarily focus on the true token $x_{t}$ and its competitive token $v^{\star}$, without excessively boosting irrelevant tokens in large vocabularies. For example, in the case where unlearning has not yet been achieved (i.e., when $p_{\theta}(x_{t}|x_{<t})>p_{\theta}(v^{\star}|x_{<t})$), the absolute value of the gradient for the true token $x_{t}$ is equal to or greater than that of $v^{\star}$ (with opposite sign) and exceeds that of other tokens (i.e., $v\neq x_{t}\ \text{and}\ v\neq v^{\star}$). This ensures efficient and targeted unlearning while avoiding the unnecessary spread of gradients across irrelevant tokens. Second, IHL resolves the issue of unbounded loss by defining a bounded loss function, ensuring that the prediction scores for tokens other than $x_{t}$ and $v^{\star}$ decrease once unlearning is complete (i.e., when $p_{\theta}(x_{t}|x_{<t})$ becomes less than $p_{\theta}(v^{\star}|x_{<t})$). This bounded loss prevents instability and divergence during unlearning. Lastly, IHL prevents the degradation of generative performance by adapting gradient updates based on the prediction scores of $x_{t}$ and $v^{\star}$, alleviating redundant updates for other tokens (i.e., $v\neq x_{t}\ \text{and}\ v\neq v^{\star}$). Compared to GA, this approach reduces redundant updates for other tokens in each sequence of the forget set $\mathcal{D}_{f}$, thereby preserving the model’s generative capabilities while achieving effective and stable unlearning.